# A Unified Framework for Forward and Inverse Problems in Subsurface Imaging using Latent Space Translations

### v1

@@ E.2 Additional Forward Modeling Results
Similar to the inverse problem, we provide detailed comparison of our models with other baseline models in Table 9. Our proposed models consistently outperform baseline models on multiple datasets indicating superior in-distributions generalizability.
@@
Metric Model FVA FVB FFA FFB CVA CVB CFA CFB STA STB
MAE $\downarrow$ FNO 0.0077 0.0385 0.0700 0.1405 0.0282 0.0829 0.0605 0.1075 0.0292 0.0379
Auto-Linear 0.0340 0.0906 0.0744 0.1537 0.0541 0.1048 0.0703 0.1356 0.0492 0.0584
Latent U-Net 0.0389 0.0473 0.0581 0.1098 0.0512 0.059 0.0576 0.0891 0.0269 0.0354
Invertible X-Net 0.0257 0.0527 0.0532 0.0887 0.0457 0.0477 0.0404 0.0671 0.0235 0.0341
MSE $\downarrow$ FNO 0.0004 0.0066 0.0259 0.0825 0.0059 0.0358 0.0251 0.0471 0.0041 0.0053
Auto-Linear 0.0084 0.0427 0.0303 0.1045 0.0217 0.0625 0.033 0.0866 0.0174 0.0173
Latent U-Net 0.0173 0.0127 0.0160 0.0489 0.0344 0.0175 0.0442 0.0297 0.0053 0.0082
Invertible X-Net 0.0062 0.0176 0.0175 0.0305 0.0185 0.0117 0.0126 0.0163 0.004 0.0087
SSIM $\uparrow$ FNO 0.9967 0.9781 0.8881 0.8354 0.9667 0.8702 0.9166 0.8160 0.9655 0.9417
Auto-Linear 0.8274 0.6081 0.5971 0.4262 0.7661 0.5808 0.6898 0.4667 0.6922 0.6631
Latent U-Net 0.9764 0.9779 0.9237 0.89 0.9659 0.9283 0.9571 0.8636 0.9819 0.9702
Invertible X-Net 0.9887 0.9782 0.9559 0.9221 0.9744 0.954 0.9757 0.9156 0.989 0.9805
@@
SSIM $\uparrow$ FNO 0.4728 0.4242 0.5690 0.5475
Auto-Linear 0.2235 0.2311 0.2569 0.2748
Latnet U-Net 0.4866 0.4567 0.5935 0.5787
Invertible X-Net 0.4408 0.4056 0.5719 0.5243
@@ E.3 Importance of Combined Loss Function for Invertible-XNet
In Figures 11 and 12, we illustrate asymmetry in learning the translation for the forward and inverse problems using CVB and CFA datasets. From the figures, we observe that the model learns the inverse mapping in the initial epochs and gradually starts to learn the solution to the forward problem in later epochs. Since the network optimizes the combined loss function on both velocity and waveform together, the gradients from the combined loss help the model to achieve better forward solution. This corroborates with our hypothesis that the model trained on combined loss is able to learn the connection between forward and inverse problem.
@@ E.4 Zero-shot Performance
E.5 Inverse Problem:
Figure 13 and 14 show generalization performance of Invertible X-Net and Latent U-Net (Large) models respectively on the inverse problem using MAE, MSE, and SSIM metrics. As described in the main paper, we show models trained across all dataset as rows and evaluated across all test datasets as columns. For comparison, we evaluate metrics such as MAE, MSE, and SSIM and calculate its difference between our models and other baseline models. For MAE and MSE, when the color intensity is blue, our model show better generalizability and vice-versa whereas, for SSIM, when the color intensity is red indicates better generalization of our model and vice-versa.
@@ E.5.1 Forward Problem:
Figures 15 and 16 compares the performance of our models Invertible X-Net and Latent U-Net (Large) against baselines. In Figure 15, we observe strong generalization of Invertible X-Net over all baselines - AutoLinear, FNO, and WaveformNet. Figure 15 (d) shows the comparative performance of Invertible X-Net against Latent U-Net (Large) where we see that Invertible X-Net dominates overall across all metrics. In Figure 16, we compare the Latent U-Net (Large) model against same baselines as above. The figure indicates Latent U-Net (Large) model has much stronger generalizability than baseline models consistently.
@@ F.1 Effect of varying Latent Space Sizes
Further, we provide visualizations of the first two primary PCA and t-SNE components of the velocity latent space in Figure 18 and 19 respectively. We take the encoder trained on a dataset and get the latent space encoding on several datasets. This visualization shows the in-distribution and out-of-distribution generality of our encoders and highlights the importance of the manifold for latent space translation. Overall, we observe that the larger latent space $70\times 70$ have better structure than $8\times 8$, indicating that the ideal size of latent space should be decided based on the complexity of dataset and problem being solved.
@@ F.3 Effect of Manifold Learning
Figure 21: Comparison of Latent U-Net Small’s performance when learning translation directly versus after reconstruction, across different training fractions.
Appendix G Additional Visualizations
Here, we provide additional visualization of the waveform and velocity predictions for baselines and our models, namely Latent-UNet (small), Latent-UNet (large), Invertible-XNet, and Invertible-XNet (cycle), for both forward and inverse problems. Please note that we show the prediction of seismic waveform and velocity in the original space by unnormalizing the predictions for every model.

### v3

@@ E.2 Additional Forward Modeling Results
Similar to the inverse problem, we provide detailed comparison of our models with other baseline models in Table 9 and 10. Our proposed models consistently outperform baseline models on multiple datasets indicating superior in-distributions generalizability.
@@
Metric Model FVA FVB CVA CVB FFA FFB CFA CFB STA STB Average Rank
MAE $\downarrow$ FNO 0.0077 0.0385 0.0700 0.1405 0.0282 0.0829 0.0605 0.1075 0.0292 0.0379 2.4
Auto-Linear 0.0340 0.0906 0.0744 0.1537 0.0541 0.1048 0.0703 0.1356 0.0492 0.0584 3.9
Latent U-Net 0.0389 0.0473 0.0581 0.1098 0.0512 0.059 0.0576 0.0891 0.0269 0.0354 2.3
Invertible X-Net 0.0257 0.0527 0.0532 0.0887 0.0457 0.0477 0.0404 0.0671 0.0235 0.0341 1.4
MSE $\downarrow$ FNO 0.0004 0.0066 0.0259 0.0825 0.0059 0.0358 0.0251 0.0471 0.0041 0.0053 1.9
Auto-Linear 0.0084 0.0427 0.0303 0.1045 0.0217 0.0625 0.033 0.0866 0.0174 0.0173 3.7
Latent U-Net 0.0173 0.0127 0.0160 0.0489 0.0344 0.0175 0.0442 0.0297 0.0053 0.0082 2.6
Invertible X-Net 0.0062 0.0176 0.0175 0.0305 0.0185 0.0117 0.0126 0.0163 0.004 0.0087 1.8
SSIM $\uparrow$ FNO 0.9967 0.9781 0.8881 0.8354 0.9667 0.8702 0.9166 0.8160 0.9655 0.9417 3.1
Auto-Linear 0.9694 0.9289 0.9038 0.8567 0.9470 0.8736 0.9253 0.8188 0.9569 0.9300 3.5
Latent U-Net 0.9764 0.9779 0.9237 0.89 0.9659 0.9283 0.9571 0.8636 0.9819 0.9702 2.3
Invertible X-Net 0.9887 0.9782 0.9559 0.9221 0.9744 0.954 0.9757 0.9156 0.989 0.9805 1.1
@@
SSIM $\uparrow$ FNO 0.8148 0.7404 0.9021 0.8447
Auto-Linear 0.6344 0.672 0.6670 0.7192
Latnet U-Net 0.8343 0.7700 0.9143 0.8626
Invertible X-Net 0.827 0.7863 0.9112 0.8500
@@ E.3 Importance of Combined Loss Function for Invertible-XNet
In Figures 12 and 13, we illustrate asymmetry in learning the translation for the forward and inverse problems using CVB and CFA datasets. From the figures, we observe that the model learns the inverse mapping in the initial epochs and gradually starts to learn the solution to the forward problem in later epochs. Since the network optimizes the combined loss function on both velocity and waveform together, the gradients from the combined loss help the model to achieve better forward solution. This corroborates with our hypothesis that the model trained on combined loss is able to learn the connection between forward and inverse problem.
@@ E.4 Zero-shot Performance
E.5 Inverse Problem
Figure 14 and 15 show generalization performance of Invertible X-Net and Latent U-Net (Large) models respectively on the inverse problem using MAE, MSE, and SSIM metrics. As described in the main paper, we show models trained across all dataset as rows and evaluated across all test datasets as columns. For comparison, we evaluate metrics such as MAE, MSE, and SSIM and calculate its difference between our models and other baseline models. For MAE and MSE, when the color intensity is blue, our model show better generalizability and vice-versa whereas, for SSIM, when the color intensity is red indicates better generalization of our model and vice-versa.
@@ E.6 Forward Problem
Figures 16 and 17 compares the performance of our models Invertible X-Net and Latent U-Net (Large) against baselines. In Figure 16, we observe strong generalization of Invertible X-Net over all baselines - AutoLinear, FNO, and WaveformNet. Figure 16 (d) shows the comparative performance of Invertible X-Net against Latent U-Net (Large) where we see that Invertible X-Net dominates overall across all metrics. In Figure 17, we compare the Latent U-Net (Large) model against same baselines as above. The figure indicates Latent U-Net (Large) model has much stronger generalizability than baseline models consistently.
@@ F.1 Effect of varying Latent Space Sizes
Further, we provide visualizations of the first two primary PCA and t-SNE components of the velocity latent space in Figure 19 and 20 respectively. We take the encoder trained on a dataset and get the latent space encoding on several datasets. This visualization shows the in-distribution and out-of-distribution generality of our encoders and highlights the importance of the manifold for latent space translation. Overall, we observe that the larger latent space $70\times 70$ have better structure than $8\times 8$, indicating that the ideal size of latent space should be decided based on the complexity of dataset and problem being solved.
@@ F.3 Effect of Manifold Learning
(a) Latent U-Net Small
(b) Invetible X-Net
Figure 22: Comparison of Latent U-Net’s and Invertible X-Net’s performance across three learning objectives: translation directly, reconstruction followed by translation, and combined learning of both, evaluated at different training fractions.
Appendix G Noise Experiments
Table 11: Quantitative results on CurveFault-B with Gaussian noise of varying variance $\sigma^{2}$ added during testing for the inverse problem.
[…]
PSNR=84.07dB PSNR=77..08dB PSNR=74.07dB PSNR=67.08dB
MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM
Latnet U-Net 185.22 93248.64 0.6930 185.54 93538.78 0.6926 187.03 94674.546 0.6917 188.56 95777.95 0.6906 195.94 101444.7 0.6844
Degradation (%) (-) (-) (-) 0.17% 0.31% 0.05% 0.97% 1.52% 0.20% 1.80% 2.71% 0.35% 5.78% 8.78% 1.24%
Invertible X-Net 180.97 86465.25 0.7117 181.32 86673.38 0.7115 182.83 87674.14 0.7106 184.17 88585.86 0.7097 192.22 94728.94 0.7036
Degradation (%) (-) (-) (-) 0.19% 0.24% 0.02% 1.02% 1.39% 0.15% 1.76% 2.45% 0.27% 6.21% 9.55% 1.13%
Auto-Linear 268.47 154713.10 0.5695 270.29 156201.81 0.5682 277.05 162156.50 0.5632 285.34 169892.26 0.5572 325.02 211616.29 0.5190
Degradation (%) (-) (-) (-) 0.67% 0.96% 0.21% 3.19% 4.81% 1.10% 6.28% 9.81% 2.15% 21.06% 36.77% 8.86%
Table 12: Quantitative results on FlatFault-B with Gaussian noise of varying variance $\sigma^{2}$ added during testing for the inverse problem.
[…]
PSNR=84.07dB PSNR=77..08dB PSNR=74.07dB PSNR=67.08dB
MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM
Latnet U-Net 92.93 33925.46 0.8515 93.47 34137.49 0.8511 95.43 34907.52 0.8499 98.19 36002.58 0.8482 108.83 40770.50 0.8417
Degradation (%) (-) (-) (-) 0.58% 0.62% 0.05% 2.68% 2.89% 0.18% 5.66% 6.12% 0.38% 17.10% 20.17% 1.15%
Invertible X-Net 94.05 32548.49 0.8532 96.23 33283.08 0.8529 104.75 36522.60 0.85 110.30 39016.04 0.8474 122.66 45509.88 0.8382
Degradation (%) (-) (-) (-) 2.31% 2.25% 0.03% 11.37% 12.20% 0.37% 17.27% 19.87% 0.67% 30.41% 39.82% 1.76%
Auto-Linear 181.39 81789.80 0.6865 187.40 85327.79 0.6821 206.58 98616.76 0.6664 221.79 110396.50 0.6511 283.50 167482.73 0.5796
Degradation (%) (-) (-) (-) 3.31% 4.32% 0.64% 13.88% 20.57% 2.93% 22.26% 34.97% 5.16% 56.28% 104.77% 15.57%
Table 13: Quantitative results on CurveFault-B with Gaussian noise of varying variance $\sigma^{2}$ added during testing for the forward problem.
[…]
PSNR=56.02dB PSNR=49.03dB PSNR=46.02dB PSNR=39.03dB
MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM
Latnet U-Net 0.0891 0.0297 0.8636 0.0891 0.0297 0.8635 0.0892 0.0297 0.8634 0.0897 0.0301 0.8630 0.0966 0.0396 0.8565
Degradation (%) (-) (-) (-) 0% 0% 0.01% 0.13% 0.30% 0.01% 0.60% 1.40% 0.06% 8.37% 33.47% 0.82%
Invertible X-Net 0.0671 0.0163 0.9157 0.0672 0.0164 0.9156 0.0674 0.0164 0.9152 0.0678 0.0166 0.9147 0.0719 0.0193 0.9090
Degradation (%) (-) (-) (-) 0.07% 0.07% 0.009% 0.42% 0.54% 0.04% 1.00% 1.68% 0.10% 7.08% 18.18% 0.73%
Auto-Linear 0.1356 0.0866 0.8188 0.1356 0.0868 0.8187 0.1360 0.0873 0.8183 0.1365 0.088 0.8179 0.139 0.0919 0.8159
Degradation (%) (-) (-) (-) 0.06% 0.14% 0.01% 0.34% 0.78% 0.05% 0.70% 1.61% 0.09% 2.53% 6.03% 0.35%
Table 14: Quantitative results on FlatFault-B with Gaussian noise of varying variance $\sigma^{2}$ added during testing for the forward problem.
[…]
PSNR=56.02dB PSNR=49.03dB PSNR=46.02dB PSNR=39.03dB
MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM MAE MSE SSIM
Latent U-Net 0.0594 0.0175 0.9283 0.0594 0.0175 0.9283 0.0598 0.0179 0.9279 0.0606 0.0187 0.9271 0.0694 0.0325 0.9178
Degradation (%) (-) (-) (-) 0% 0% 0% 0.66% 2.16% 0.04% 1.99% 6.99% 0.12% 16.87% 85.99% 1.12%
Invertible X-Net 0.0477 0.0117 0.9540 0.0478 0.0117 0.9540 0.0482 0.0119 0.9536 0.0487 0.0122 0.9531 0.0535 0.016 0.9463
Degradation (%) (-) (-) (-) 0.17% 0.33% 0% 0.94% 1.88% 0.04% 1.93% 4.26% 0.10% 12.16% 36.39% 0.80%
Auto-Linear 0.1048 0.0625 0.8736 0.1051 0.0628 0.8734 0.1061 0.0641 0.8726 0.1072 0.0656 0.8717 0.112 0.0727 0.868
Degradation (%) (-) (-) (-) 0.25% 0.54% 0.02% 1.19% 2.59% 0.11% 2.24% 5.06% 0.21% 6.86% 16.37% 0.64%
Appendix H Additional Visualizations
Here, we provide additional visualization of the waveform and velocity predictions for baselines and our models, namely Latent-UNet (small), Latent-UNet (large), Invertible-XNet, and Invertible-XNet (cycle), for both forward and inverse problems. Please note that we show the prediction of seismic waveform and velocity in the original space by unnormalizing the predictions for every model (Figures 23 - 34).
@@ Appendix H Additional Visualizations
Appendix I Code Appendix
All the code required to train and evaluate the proposed methods, as well as the baselines, has been uploaded to an anonymous GitHub repository: https://github.com/KGML-lab/Generalized-Forward-Inverse-Framework-for-DL4SI. The data and corresponding processing code used in this work are sourced from the OpenFWI website: https://openfwi-lanl.github.io/docs/data.html.